January 5, 2011

Observers: Jill Bechtold, Dave Thompson

Completely overcast all afternoon and at sunset. Not looking good for tonight!

LUCIFER's mask moving mechanism is malfunctioning and it has been decided to not allow any spectroscopic observations for the rest of the AZ run.

LBC-Blue's filter wheel is malfunctioning. Day crew had LBC-Blue apart this afternoon, but could not figure out how to get to the right part of it to fix the filter wheel encoder, which was one theory on why it wasn't working. The other theory has to do with cycling the counter up to 6 million steps. So they put it back together and Dave and Olga are going to recalibrate the filter wheel, and if the problem is the 6 million step thing, it should work.

At sunset, it was proving difficult to get LBC-Blue and Red to come back on, and there is something wrong with the Red cryostat.

8pm. Momentary sucker hole, but we decided to keep working on LBC-Blue and it disappeared.

9:20pm. Olga thinks she has recalibrated LBC-Blue's filter wheels. Still extremely cloudy, so we're going to try one of Peter Milne's SNe with LBC. But there are more problems again with LBC port errors. Switching from Lucifer to LBC and will open soon. LBC is having troubles powering up, Olga has to try it numerous times for it to come on.

9:47pm. LBC is on, at last. Opening the dome. Very cloudy. Olga will cycle through the filters and look at sky counts to check that everything is OK.

9:56pm. LBC crashed again.

10:38pm Went to SN09ig but it was far over and couldn't do dofpia for LBC because it was too cloudy. So we went to SN05mz, which is near zenith now, doubled the exposure but still couldn't collimate. Clouds are worse. So we closed. Olga went home. THANKS FOR YOUR HELP OLGA!

4:00am Still completely cloudy. Maybe tomorrow will be better.
Topic revision: r5 - 12 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback